# Behavior-Regularized Diffusion Policy Optimization for Offline Reinforcement Learning

### v1

@@ 5 Experiments
5.1 Results of Synthetic 2D Datasets
We leverage synthetic 2D datasets used in Lu et al. (2023) to further our understanding of the mechanism of BDPO. Each dataset contains data points $x_{i}$ paired with specific energy values $Q(x_{i})$. By re-sampling the data according to the Boltzmann distribution $p(x_{i})\propto\exp(Q(x_{i})/\eta)$, we can obtain the ground truth target distribution at a specific regularization strength $\eta$ (Haarnoja et al., 2018; Nair et al., 2020). As an example, Figure 5 illustrates both the original dataset and the re-sampled dataset with $\eta=0.06$ for the 8gaussians and 2spirals datasets.
Our objective is to analyze the properties of the diffusion policy and the diffusion value function trained with BDPO. The results are depicted in Figure 4 (full results on all datasets in Figure 11), where each row corresponds to a different dataset and visualizes the iterative sampling process. From left to right, the samples evolve from an initial noisy distribution toward a well-defined structure as the reverse diffusion progresses. In the initial steps ($n=50$ to $n=30$), the sample movement is subtle, while in the later steps ($n=30$ to $n=0$), the samples rapidly converge to the nearest modes of the data. The final column on the right presents the ultimate samples generated by the policy, which closely match the ground-truth target distribution visualized in Figure 5.
(a) Sensitivity analysis of $\eta$.
(b) Sensitivity analysis of $\rho$.
Figure 7: Sensitivity analysis of (a) the regularization strength $\eta$ and (b) the lower confidence bound coefficient $\rho$ of BDPO. For each configuration, we report the mean and the standard deviation of the performances aggregated from 5 independent seeds and 10 evaluation episodes for each seed.
Besides the generation result, the background color in Figure 4 depicts the landscapes of the diffusion value functions $V^{\pi}$ at $n=50/30/15/5/0$. In earlier, noisier steps (e.g., $n=50$), the output values across the space are similar, resulting in a smoother landscape and weaker guidance during sampling. As the noise level decreases in later steps (e.g., $n=5$), the value outputs vary more sharply, creating stronger guidance on the diffusion policy. Such progression enables the model to refine samples while exploring sufficiently across the space, finally yielding results that align with the target distribution.
5.2 Results on Continuous Control Benchmark
To assess the performance of BDPO on complex continuous control problems, we utilize datasets from D4RL as our test bench and compare BDPO against a wide spectrum of offline RL algorithms. Specifically, we include 1) CQL (Kumar et al., 2020b) and IQL (Kostrikov et al., 2022), which are representatives using straightforward one-step policies; 2) Decision Diffuser (DD) (Ajay et al., 2023), which uses diffusion models for planning and extracts actions from planned trajectories; 3) SfBC (Chen et al., 2023), IDQL (Hansen-Estruch et al., 2023), and QGPO (Lu et al., 2023), which use diffusion to model the behavior and further refine actions using the value functions; 4) Diffusion-QL (Wang et al., 2023), which propagates the gradient of Q-value functions over the whole generation process; and 5) DAC (Fang et al., 2024), which aligns the output of diffusion policies with the gradient of Q-values.
The results are listed in Table 1. Our findings indicate that diffusion-based methods, particularly those with diffusion-based actor and regularization (including BDPO, DAC and Diffusion-QL), substantially outperform their non-diffusion counterparts especially in locomotion tasks. Meanwhile, BDPO consistently achieves superior performance across nearly all datasets, underscoring the effectiveness of combining diffusion policies and behavior-regularized RL framework. Besides the final performance, we plot the training curves of BDPO in Section E.3. Overall, BDPO achieves fast and stable convergence except for some of the antmaze datasets, where variations may occur due to the sparse reward nature of these datasets.
@@ 5.3 Analysis of BDPO
Policy Parameterization. As discussed in Section 1, improper assumptions about the action distribution may lead to inaccurate penalty calculations and ultimately degrade the overall performance. To validate this claim, we conduct a series of ablation studies that replace the actor parameterization with Gaussian distributions and Dirac distributions, respectively, while keeping other experimental configurations unchanged. For a fair comparison, we fine-tune the regularization strength $\eta$ by searching within a predefined range and report the performance corresponding to the optimal value. The results are illustrated in Figure 6, with additional details regarding the implementations provided in Appendix B.3. While simpler architectures generally converge more quickly, their performance at convergence often lags behind that of BDPO, indicating a limitation in their capacity in policy parameterization.
Effect of Regularization Strength $\eta$. The hyper-parameter $\eta$ controls the regularization strength: larger $\eta$ forces the diffusion policy to stay closer to the behavior diffusion, while smaller $\eta$ places more emphasis on the guidance from value networks. Figure 7(a) demonstrates this effect: In general, smaller $\eta$ does lead to better performances. However, excessively small values can result in performance degradation, as observed in walker2d-medium-v2. In contrast to auto-tuning $\eta$ via dual gradient descend (Fang et al., 2024; Zhang et al., 2024a), we find that adjustable $\eta$ causes fluctuations in penalty calculation. Therefore, we employ a fixed $\eta$ throughout the training.
Effect of Lower Confidence Bound Coefficient $\rho$. The hyper-parameter $\rho$ determines the level of pessimism on OOD actions since the variances of Q-values on these actions are comparatively higher than in-dataset actions. In Figure 7(b), we discover that there exists a certain range of $\rho$ where the lower confidence bound value target works best, while excessively larger or smaller values lead to either underestimation or overestimation in value estimation.
@@ 6 Conclusions
Our core contribution is to extend the behavior-regularized RL framework to diffusion policies by implementing the regularization as the accumulated discrepancies of each single-step transition of the reverse diffusion. Theoretically, we demonstrate the equivalency of the pathwise KL and the commonly adopted KL constraint on action distributions. In practice, we instantiate the framework with an actor-critic style algorithm which leverages value functions in two time scales for efficient optimization. The proposed algorithm, BDPO, produces generation results that closely align with the target distribution, thereby enhancing its performance and applicability in complex continuous control problems.
@@ Appendix A Introduction about the Benchmarks
Synthetic 2D Datasets. We leverage the synthetic 2D datasets from Lu et al. (2023) as a sanity check and a convenient illustration of the mechanism of BDPO. This task set encompasses 6 different datasets. For each dataset, the data points $x_{i}$ are associated with different energies $e_{i}$. By sampling from the Boltzmann distribution associated with the temperature $\eta$:
$p(x_{i})\propto\exp(e_{i}/\eta),x_{i}\in\mathcal{D},$
we can obtain ground truth sample results regularization strength $\eta$ (see Figure 8).
Figure 8: Illuration of synthetic 2D datasets and sampling results with various temperature $\eta$.
@@
Appendix B Hyper-parameter Configurations
B.1 D4RL Datasets
We largely follow the configurations from DAC (Fang et al., 2024). Parameters that are common across all datasets are listed in Table 2. The hyper-parameters of value networks in the table are the same for both the Q-value networks $Q_{\psi}$ and diffusion value networks $V_{\phi}.$
Table 2: Common hyper-parameters across all datasets.
Diffusion Backbone MLP(512, 512, 512, 512)
@@ B.1 D4RL Datasets
Actor Learning Rate 0.00001
@@
Actor EMA Rate 0.005
Actor Samples 10
@@
Pretrain Steps 2000000
Train Steps 2000000
Value Warmup Steps 50000
@@
halfcheetah-medium-v2 0.5 0.5 False
@@
hopper-medium-v2 2.0 2.0 False
@@
walker2d-medium-v2 1.0 1.0 False
@@
antmaze-umaze-v2 1.0 5.0 True
antmaze-medium-play-v0 1.0
antmaze-large-play-v0 10.0
antmaze-umaze-diverse-v0 0.5 5.0 True
antmaze-medium-diverse-v0 1.0
antmaze-large-diverse-v0 5.0
@@ B.3 Details about the Ablation on Policy Parameterization
In this section, we provide more details about how we implement the variants that use deterministic policies or Gaussian policies. The only difference between these variants and BDPO lines in the actor parameterization, which means that we keep the implementation of the value networks (e.g. ensemble tricks and LCB value target) and the hyper-parameters (e.g. the LCB coefficient $\rho$) identical to BDPO. We inherit the general behavior-regularized RL framework defined in Eq. (1), where we update the soft Q-value functions using the target
@@
$\max_{\pi}\ -\eta D(\pi,\nu)+\gamma\mathbb{E}_{a\sim\pi}\left[Q^{\pi}(s,a)\right],$ (20)
@@
For Gaussian policies, we parameterize it as Diagonal $\pi(\cdot|s)=\mathcal{N}(\mu_{\pi}(s),\sigma_{\pi}^{2}(s))$, and the regularization is implemented as the KL divergence:
$D(\pi,\nu)=\operatorname{KL}\left[\pi\|\nu\right]=\log\frac{\sigma_{\nu}}{\sigma_{\pi}}+\frac{\sigma_{\pi}^{2}+(\mu_{\pi}-\mu_{\nu})^{2}}{2\sigma_{\nu}^{2}}-\frac{1}{2}.$ (22)
@@
halfcheetah-medium-expert-v2 10 0.75
walker2d-medium-v2 1.5 0.25
@@ Appendix C Proof for the Main Results
$\mathbb{E}_{a^{0}\sim\nu(s),a^{n}\sim q_{n|0}(\cdot|a^{0})}\left[\operatorname{KL}\left[q_{n-1|0,n}(\cdot|a^{0},a^{n})\|p^{\nu,s}_{n-1|n}(\cdot|a^{n})\right]\right]$ (23)
for $n\in\{0,1,\ldots,N\}$. The optimal solution $p_{n-1|n}^{*,s}$ of Eq. (13) is the reverse process corresponding to the forward process $q_{0:N}^{\pi^{*},s}=\pi^{*}(a^{0}|s)\prod_{n=1}^{N}q_{n|n-1}(a^{n}|a^{n-1})$, where $\pi^{*}(a|s)\propto\nu(a|s)\exp(Q(s,a)/\eta)$.
@@
First, we note that the Markov process $p_{0:N}^{\nu,s}$ is the reverse process corresponding to forward diffusion process $q^{\nu,s}(u_{0:N})=\nu(a^{0}|s)\prod_{n=1}^{N}q(a^{n}|a^{n-1})$, i.e.,
$p^{\nu,s}_{n-1|n}=q^{\nu,s}_{n-1|n}$ (24)
where
$q^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})\coloneqq\frac{\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}[q_{n-1|0}(a^{n-1}|a^{0})q_{n|n-1}(a^{n}|a^{n-1})]}{\mathbb{E}_{a^{0}\sim\nu(s)}[q_{n|0}(a^{n}|a^{0})]}.$ (25)
This follows from the fact that $p_{n-1|n}^{\nu,s}$ minimizes the objective in Eq. (23)). Solving the constrained optimization problem
$\min_{p_{n-1|n}^{\nu,s}(\cdot|a^{n})}\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})\operatorname{KL}\left[q_{n-1|0,n}(\cdot|a^{0},a^{n})\|p^{\nu,s}_{n-1|n}(\cdot|a^{n})\right]\right]$

### v2

@@ 5.3 Analysis of BDPO
Regularization Strength $\eta$. The hyperparameter $\eta$ controls the regularization strength: larger $\eta$ forces the diffusion policy to stay closer to the behavior diffusion, while smaller $\eta$ places more emphasis on the guidance from value networks. Figure 6 demonstrates this effect: in general, smaller $\eta$ does lead to better performances. However, excessively small values can result in performance degradation, as observed in walker2d-medium-v2. In contrast to auto-tuning $\eta$ via dual gradient descend (Fang et al., 2024; Zhang et al., 2024a), we find that adjustable $\eta$ causes fluctuations in penalty calculation. Therefore, we employ a fixed $\eta$ throughout training.
Lower Confidence Bound Coefficient $\rho$. The hyperparameter $\rho$ determines the level of pessimism on OOD actions since the variances of Q-values on these actions are comparatively higher than in-dataset actions. In Figure 7, we discover that there exists a certain range of $\rho$ where the lower confidence bound value target works best, while excessively larger or smaller values lead to either underestimation or overestimation in value estimation.
Table 2: Runtime comparison among BDPO, DAC, and Diffusion-QL (denoted as D-QL in the table), measured in minutes.
Alg Phase $N=5$ $N=20$ $N=50$
BDPO policy 17 20 18
value 63 135 285
total 80 155 303
DAC policy 30 31 30
value 30 109 251
total 60 140 285
D-QL policy 57 132 308
value 31 80 252
total 88 212 560
Policy Parameterization. As discussed in Section 1, improper assumptions about the action distribution may lead to inaccurate penalty calculations and ultimately degrade the overall performance. To validate this claim, we conduct a series of ablation studies that replace the actor parameterization with Gaussian distributions and Dirac distributions, respectively, while keeping other experimental configurations unchanged. For a fair comparison, we fine-tune the regularization strength $\eta$ by searching within a predefined range and report the performance corresponding to the optimal value. The results are illustrated in Figure 8, with additional details regarding the implementations provided in Appendix B.3. While simpler architectures generally converge more quickly, their performance at convergence often lags behind that of BDPO, indicating a limitation in their capacity for policy parameterization.
Discussion on Runtime. Table 2 compares the runtime of BDPO, DAC, and Diffusion-QL using the walker2d-medium-v2 dataset. Our method consists of three key steps: pretraining behavior diffusion, training value functions ($Q^{\pi}$ and $V^{\pi,s}_{n}$), and optimizing the actor. The pertaining phase takes around 8 minutes and is therefore omitted in further analysis. The $Q^{\pi}$ training follows the same approach as Diffusion-QL and DAC, involving diffusion path sampling at the next state, while $V^{\pi,s}_{n}$ introduces acceptable overhead since it only requires single-step diffusion. For actor training, both DAC and BDPO use single-step diffusion, ensuring computational costs remain constant regardless of diffusion steps. In contrast, Diffusion-QL needs to back-propagate through the entire diffusion path, leading to a significantly higher runtime, especially when the number of diffusion steps increases.
@@ 6 Conclusions
Our core contribution is to extend the behavior-regularized RL framework to diffusion policies by implementing the regularization as the accumulated discrepancies of each single-step transition of the reverse diffusion. Theoretically, we demonstrate the equivalence of the pathwise KL and the commonly adopted KL constraint on action distributions. In practice, we instantiate the framework with an actor-critic style algorithm that leverages value functions on two time scales for efficient optimization. The proposed algorithm, BDPO, produces generation results that closely align with the target distribution, thereby enhancing its performance and applicability in complex continuous control problems.
Acknowledgements
This work is primarily supported by the National Science Foundation of China (62276126, 62250069). Chenjun Xiao gratefully acknowledges funding from the National Natural Science Foundation of China (62406271).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Introduction about the Benchmarks
Synthetic 2D Datasets. We leverage the synthetic 2D datasets from Lu et al. (2023) as a sanity check and a convenient illustration of the mechanism of BDPO. This task set encompasses 6 different datasets. For each dataset, the data points $x_{i}$ are associated with different energies $\mathcal{E}(x_{i})$. By sampling from the Boltzmann distribution associated with the temperature $\eta$:
$p(x_{i})\propto\exp(\mathcal{E}(x_{i})/\eta),x_{i}\in\mathcal{D},$
we can obtain ground-truth sample results at a specific regularization strength $\eta$ (see Figure 9).
Figure 9: Illustration of synthetic 2D datasets and sampling results with various temperature $\eta$.
@@
Metrics. We measure the performance using the normalized score, which can be calculated by
[…]
where $\text{Score}(\pi)$ is the average cumulated rewards achieve by the policy $\pi$, $\text{Score}_{\text{random}}$and $\text{Score}_{\text{expert}}$ are reference scores for random policies and expert policies, provided by D4RL.
Appendix B Hyper-parameter Configurations
B.1 D4RL Datasets
We largely follow the configurations from DAC (Fang et al., 2024). Parameters that are common across all datasets are listed in Table 3. The hyperparameters of value networks in the table are the same for both the Q-value networks $Q_{\psi}$ and diffusion value networks $V_{\phi}.$
@@ B.1 D4RL Datasets
Table 3: Common hyperparameters across all datasets.
Diffusion Backbone MLP(256, 256, 256, 256) for locomotion tasks MLP(512, 512, 512, 512) for antmaze tasks
@@
Behavior Learning Rate 0.0003 for locomotion tasks and 0.0001 for antmaze tasks
Actor Learning Rate 1e-5
@@
Actor EMA Rate 5e-3
Actor Sampling For antmaze-umaze-diverse-v0, generate 1 sample and execute For others, generate 10 samples and select the action with the highest Q
@@
Pretrain Steps 2e6
Train Steps 2e6
Value Warmup Steps 5e4
@@
halfcheetah-medium-v2 0.5 0.05 False
@@
hopper-medium-v2 2.0 0.2 False
@@
walker2d-medium-v2 1.0 0.15 False
@@
antmaze-umaze-v2 0.8 0.5 True
antmaze-umaze-diverse-v0
antmaze-medium-play-v0 0.8 0.2 True
antmaze-medium-diverse-v0
antmaze-large-play-v0 0.8 1.0 True
antmaze-large-diverse-v0
@@ B.3 Details about the Ablation on Policy Parameterization
In this section, we provide more details about how we implement the variants that use deterministic policies or Gaussian policies. The only difference between these variants and BDPO lines in the actor parameterization, which means that we keep the implementation of the value networks (e.g., ensemble tricks and LCB value target) and the hyperparameters (e.g., the LCB coefficient $\rho$) identical to BDPO. We inherit the general behavior-regularized RL framework defined in Eq. (1), where we update the soft Q-value functions using the target
@@
$\max_{\pi}\ \gamma\mathbb{E}_{a\sim\pi}\left[Q^{\pi}(s,a)\right]-\eta D(\pi,\nu),$ (21)
@@
For Gaussian policies, we parameterize it as Diagonal Gaussian distribution $\pi(\cdot|s)=\mathcal{N}(\mu_{\pi}(s),\sigma_{\pi}^{2}(s))$, and the regularization is implemented as the KL divergence:
$D(\pi(s),\nu(s))=D_{\rm KL}\left[\pi(s)\|\nu(s)\right]=\log\frac{\sigma_{\nu}(s)}{\sigma_{\pi}(s)}+\frac{\sigma_{\pi}^{2}(s)+(\mu_{\pi}(s)-\mu_{\nu}(s))^{2}}{2\sigma_{\nu}^{2}(s)}-\frac{1}{2}.$ (23)
@@
halfcheetah tasks 0.25 0.75
hopper tasks 1.0 0.1
walker2d tasks 1.5 0.25
@@ Appendix C Proof for the Main Results
$\mathbb{E}_{a^{0}\sim\nu(s),a^{n}\sim q_{n|0}(\cdot|a^{0})}\left[D_{\rm KL}\left[q_{n-1|0,n}(\cdot|a^{0},a^{n})\|p^{\nu,s}_{n-1|n}(\cdot|a^{n})\right]\right]\quad\forall n\in\{1,2,\ldots,N\},s\in\mathcal{S}.$ (24)
Then, for any given Q-value function $Q(s,a)$, the optimal solution $p^{*,s}$ of Eq. (14) is the reverse process corresponding to the forward process $q^{\pi_{Q},s}_{0:N}=\pi_{Q}(a^{0}|s)\prod_{n=1}^{N}q_{n|n-1}(a^{n}|a^{n-1})$, where $\pi_{Q}(a|s)\propto\nu(a|s)\exp(Q(s,a)/\eta)$.
@@
Let’s first consider the behavior diffusion policy $p^{\nu}$, which minimizes the following problem for every $n$ and state $s$:
$\min_{p^{\nu,s}_{n-1|n}}\quad\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})D_{\rm KL}\left[q_{n-1|0,n}(\cdot|a^{0},a^{n})\|p_{n-1|n}^{\nu,s}(\cdot|a^{n})\right]\right]$